# Efficient Multi Subject Visual Reconstruction from fMRI Using Aligned Representations

### v1

@@ Abstract
This work introduces a novel approach to fMRI-based visual image reconstruction using a subject-agnostic common representation space. We show that the brain signals of the subjects can be aligned in this common space during training to form a semantically aligned common brain. This is leveraged to demonstrate that aligning subject-specific lightweight modules to a reference subject is significantly more efficient than traditional end-to-end training methods. Our approach excels in low-data scenarios. We evaluate our methods on different datasets, demonstrating that the common space is subject and dataset agnostic.
Machine Learning, ICML
@@ 1 Introduction
Over the past several decades, the use of machine learning techniques has enabled decoding and/or reconstruction of information represented in neural activation patterns measured in awake, behaving humans. Because the neural representations supporting behavior likely entail patterns of activity across large swaths of tissue, these approaches are thought to offer a key step towards characterizing and quantifying the computational principles brains use to support cognition (Naselaris et al., 2011).
Early approaches, including self-supervised learning frameworks  (Beliy et al., 2019; Gaziv et al., 2022; Mozafari et al., 2020; Shen et al., 2019; Seeliger et al., 2018; Ren et al., 2021) struggled to generate semantically accurate image reconstructions, often resulting in images with limited resemblance to the original stimuli. Furthermore, these methods required large amounts of data for training, and their generalization to a wider population of subjects was severely limited. Significant advancements were achieved by incorporating multimodal data, such as text (Lin et al., 2022), and later by leveraging the generative power of latent diffusion models (LDMs) (Takagi & Nishimoto, 2023; Scotti et al., 2023; Lu et al., 2023; Xia et al., 2024). These models introduced novel mapping techniques to project fMRI signals into the input space of LDMs, yielding higher-quality natural images. However, despite the improved fidelity, the issue of generalization to new participants persisted. In response, (Scotti et al., 2024) proposed projecting fMRI data into a shared representation space, demonstrating significant improvements in both performance and generalizability.
Beyond being able to use decoding and reconstruction approaches to provide evidence of the existence of information represented in a particular brain activation pattern, an additional goal is to compare decoding or reconstruction performance across experimental manipulations, such as cued visual attention  (Serences & Boynton, 2007; Kamitani & Tong, 2005; Scolari et al., 2012; Sprague & Serences, 2013; Itthipuripat et al., 2019; Sprague et al., 2018) and visual working memory  (Serences et al., 2009; Harrison & Tong, 2009; Ester et al., 2013; Sprague et al., 2014; Christophel et al., 2012; Li et al., 2021) Ḣowever, for these approaches to be useful for a cognitive neuroscience lab, the lab would need to acquire sufficient data to train a decoding or reconstruction model for each individual participant (often >20 per study to attain statistical power). The state-of-the-art methods described above use extremely massive datasets with dozens of hours of fMRI data per participant, which is not feasible in almost all cases. Thus, there is a clear need to maximize the efficiency and generalizability of image reconstruction methods such that a model for a new participant can be trained using as little data as possible.
In this work, we introduce an efficient and generalizable approach to fMRI-based visual reconstruction that leverages a subject-agnostic common representation space. Unlike previous approaches that perform end-to-end training of fMRI signals to mapped contextual embeddings, we propose aligning brain activity from new subjects to a pre-trained reference subject in a shared representation space using Adapter Alignment (AA) training. Throughout this work, we refer to lightweight subject-specific modules designed to map input fMRI signals to a shared representation space as ”adapters” (Liu et al., 2024). These modules are implemented as single-layer neural networks. In their linear form, adapters consist of a single linear transformation layer, while in the non-linear form, they include a linear layer followed by a non-linear activation function such as GELU (Hendrycks & Gimpel, 2023). Using AA training, we significantly reduce the amount of fMRI data required and the training time necessary to achieve accurate reconstructions and improve the reconstruction quality. Our method achieves better performance than the training method used by existing state-of-the-art models, but with far fewer training hours, making it feasible for cognitive neuroscience labs with limited data collection resources. We build on this approach by also proposing an adaptive image selection strategy, allowing for optimal fine-tuning of new subjects with minimal data, further improving the practicality and accessibility of fMRI-based image reconstruction. Finally, we evaluate our proposed methods on different datasets.
Our contributions to this work are as follows:
  • We provide strong evidence for the existence of a common visual representation space across subjects, where brain activity from different individuals can be aligned in a shared space. By extracting and analyzing embeddings from a pre-trained model, we show that even without explicit alignment, subject-specific fMRI signals, when mapped to a common space display similar behavior and representation structures. Furthermore, we demonstrate that applying a simple orthogonal transformation improves the alignment of embeddings between subjects, and using non-linear adapters with explicit mapping leads to almost perfect alignment across subjects in the common space. (Section 2)
  • We introduce Adapter Alignment (AA), a novel approach to facilitate efficient subject-specific alignment in fMRI-based visual reconstruction. Our training pipeline first pre-trains on a reference subject, then aligns new subjects’ brain data using lightweight adapters that map them to the shared common space. This process, which is more efficient than traditional end-to-end training, enables faster convergence and better performance, particularly in low-data scenarios. By incorporating MSE losses at the adapter level for common images, we show that our method not only reduces training time but also improves generalization and reconstruction quality. We also present scaling experiments to show that AA training is scalable and agnostic to the components used in the pipeline. (Section 3, Appendix Section A)
  • We develop a greedy image selection algorithm that identifies the most representative subset of images for training new subjects. By efficiently selecting fMRI signals that capture the most informative aspects of the common space, this strategy significantly reduces the amount of data required for fine-tuning without sacrificing performance (from a baseline of 250 images to 150 images). This makes our method more practical for real-world applications, where acquiring large fMRI datasets is often prohibitively expensive. (Section 4)
  • We validate the generalizability of our approach by applying it to data acquired from different MRI scanners. In particular, we leverage the THINGS dataset  (Hebart et al., 2023) to fine-tune our pre-trained models, originally trained on the NSD dataset. Despite the variations in scanning protocols and data characteristics, our Adaptive Alignment (AA) strategy continues to outperform conventional fine-tuning. (Section 5)
@@ 2 A Common Visual Representation Space
Central to our approach is the hypothesis that each individual’s brain (or at least the tissue involved in processing static visual images) can be considered an instantiation of a ’common’ brain, with activity patterns varying along similar dimensions in a common shared representational space. Aspects of this hypothesis are already strongly supported: most human brains have localized tissue responsive to specific types of stimuli, like faces and scenes. Additionally, several continuous dimensions of high-level visual selectivity have been identified across the ventral temporal cortex (Huth et al., 2012), including dimensions like animate/inanimate, large/small (based on real-world size) (Long et al., 2018), and spiky/round (Bao et al., 2020). Retinotopic organization is another example of a shared representational space, as each individual brain instantiates a unique anatomical mapping from retinal coordinates to cortical representations, and adaptive models can be estimated to transform activation patterns from individual brains to this common shared retinotopic coordinate system  (Sprague & Serences, 2013; Wandell & Winawer, 2015). We analyze the similarity of representation spaces from different perspectives of common concepts, visual analysis, and eigenvectors in the following subsections. Finally, we reflect on the different ways of aligning the common representations.
@@
Existing works, such as MindEye2 (Scotti et al., 2024), employ a simple linear adapter network to project subject-specific fMRI signals into a common representation space. A key indication of the existence of such a shared space is the structural similarity of embeddings across subjects. To explore this, we first use a pre-trained MindEye2 model to extract common-space embeddings for the shared images (the test set in MindEye2) across all subjects. Initial cosine similarity between the embeddings from different subjects is low. However, after applying an orthogonal transformation using Procrustes analysis (Schönemann, 1966), the cosine similarity improves, suggesting that the embeddings can be structurally aligned. Despite this improvement, the alignment is not perfect, likely due to the limitations of linear transformations.
(a) Weak alignment before an orthogonal transformation.
(b) Improved alignment after orthogonal transformation.
(c) Near perfect alignment with nonlinear adapters.
Figure 1: Cosine similarity of shared image embeddings in the common space. (a) The alignment is weak before an orthogonal transformation. (b) Alignment improves after orthogonal transformation. (c) Near perfect alignment is achieved with nonlinear adapters and explicit alignment. Only subjects 1, 2, 5, and 7 are considered in this plot, as these subjects were used during training.
As a preliminary experiment, we pre-trained a single-subject model with a non-linear adapter and explicitly mapped subsequent subjects to this reference subject using subject-specific non-linear adapters. This approach results in near-perfect alignment in the common space. Although existing methods do not explicitly align subject-specific representations, we observed significant semantic and structural similarities across subjects’ representations. Furthermore, it is possible to map one subject’s representations to another’s using non-linear adapters. While this alignment in the common space is nearly perfect, it does not directly translate to better reconstruction performance.
We also analyze the singular vectors from the adapter weights to present semantic category analysis and images in the representation space to perform visual analysis. We present our results and analysis in the Appendix Section C. This evidence supports the core hypothesis of our work: it is possible to construct an aligned, subject-agnostic common space. In Section 3, we empirically test this hypothesis by mapping a new subject to a reference subject during training and evaluating its reconstruction performance. Additional analysis is provided in Appendix B.
3 fMRI Reconstruction
Building a common brain representation requires constructing an efficient fMRI reconstruction pipeline. Our approach is guided by two primary objectives: 1) achieving competitive reconstruction performance and 2) prioritizing efficiency over marginal performance gains. We hypothesize that it is possible to pre-train a single reference brain and align subsequent subjects to this reference using subject-specific adapters, significantly reducing the computational cost while maintaining high reconstruction quality.
@@ 3.1 Dataset
We use the Natural Scenes Dataset (NSD) (Allen et al., 2022) but apply a different split compared to the traditional setup. Typically, common images across subjects are assigned to the test set, while the training set consists of unique images viewed by each subject. However, for effective Common Space Alignment, it is critical to have a one-to-one mapping of images across subjects. To achieve this, we incorporate the common images into the training set and swap an equal amount of training images to the test sets. This means that every subject now has a unique set of images on which they are tested. The test split is chosen to roughly ensure an equal distribution across different categories in the dataset. We present the performance of the original training method and the proposed technique on the new split to ensure a fair comparison across subjects and alignment techniques.
@@ 3.2 Adapter Alignment
Figure 2: Training procedures for fMRI-Image Reconstruction. Top: Traditional Approach to reconstruction. All subjects are trained simultaneously and loss functions are applied at the MLP Mapper to align the output to pre-trained CLIP space. Bottom: Proposed AdapterAlign training pipeline. First you train a single reference subject and obtain their embeddings in the shared space. Then you train subsequent subjects to minimize loss at both the shared space and the mapper. The same mapper weights are transferred over in step 2.
@@
Current approaches for training multi-subject pipelines with a common representation space typically follow an end-to-end strategy, where fMRI inputs are processed by the model and the loss is computed at the final layer to minimize the discrepancy between the ground-truth and predicted embeddings. In this setup, the common space emerges organically during training without explicit constraints. However, this approach has two key drawbacks: (1) it requires extended training to form a satisfactory common space and achieve convergence, and (2) the subject-specific embeddings in the common space are not strongly aligned, making it harder for the residual MLP to map multiple subject inputs to a unified output.
@@
To address these challenges, we propose Adapter Alignment (AA) training. In AA, the pipeline is first pre-trained end-to-end on a single reference subject, chosen based on reconstruction performance and data availability. Among NSD subjects, Subject 1 consistently exhibited the best reconstruction performance, making it the ideal candidate for constructing the reference space. After pre-training, the output embeddings for the subject-specific adapter, particularly those corresponding to the shared images, are extracted.
@@
When fine-tuning a new subject, we adopt a two-step process. First, we train the new subject’s adapter using the shared images, with the objective of aligning its output to the embeddings of Subject 1 for the same images. The adapter is allowed to overfit on these shared images until the loss is minimized. In the second step, we unfreeze the entire pipeline and resume training, applying the usual losses at the final output. Additionally, we introduce an MSE loss at the adapter level specifically for the common images, which allows the adapter to restructure itself to account for subject-specific discrepancies and adapt to the unique images encountered by the new subject. We call this process of overfitting the adapter and fine-tuning end to end ”AAMax”.
@@
This training method offers improvements in both high-data and low-data scenarios. When all training data is available, AAMax converges faster while achieving performance similar to traditional end-to-end training. However, in limited data settings, AAMax proves significantly more efficient, yielding better reconstructions and higher accuracy with fewer training samples. We outline our method in Figure 2 and break down our pipeline in the following section.

### v2

@@ Abstract
Reconstructing visual images from fMRI data is a central but highly challenging problem in neuroscience. Despite recent progress, current methods fall short when data and computation are limited—–precisely the conditions under which this task is most critical. We introduce a novel architecture-agnostic training paradigm to improve fMRI-based visual reconstruction through a subject-agnostic common representation space. We show that it is possible to leverage subject-specific lightweight modules to develop a representation space where different subjects not only lie in a shared space but are also aligned semantically. Our results demonstrate that such a training pipeline achieves significant performance gains in low-data scenarios. We supplement this method with a novel algorithm to select the most representative subset of images for a new subject. Using both techniques together, one can fine-tune with at most 40% of the data while outperforming the baseline trained with the minimum standard dataset size. Our method generalizes across different training paradigms and architectures, producing state-of-the-art performance and demonstrating that a subject-agnostic aligned representation space is the next step towards efficient Brain-Computer Interfaces.
@@ 1 Introduction
Over the past several decades, the use of machine learning techniques has enabled the decoding and/or reconstruction of information represented in neural activation patterns. These approaches offer a key step toward characterizing and quantifying cognition Naselaris et al. (2011). Beyond establishing the presence of information in brain activity, another important goal is to compare decoding or reconstruction performance across experimental manipulations, such as cued visual attention (Serences and Boynton, 2007; Kamitani and Tong, 2005; Scolari et al., 2012; Sprague and Serences, 2013; Itthipuripat et al., 2019; Sprague et al., 2018) and visual working memory (Serences et al., 2009; Harrison and Tong, 2009; Ester et al., 2013; Sprague et al., 2014; Christophel et al., 2012; Li et al., 2021). However, for these approaches to be practical in a cognitive neuroscience laboratory, it is typically necessary to acquire large amounts of data.
Recent efforts have begun to address the challenges associated with decoding and reconstructing brain activity. Early approaches, including self-supervised learning frameworks  (Beliy et al., 2019; Gaziv et al., 2022; Mozafari et al., 2020; Shen et al., 2019; Seeliger et al., 2018; Ren et al., 2021) struggled to generate semantically accurate image reconstructions, often resulting in images with limited resemblance to the original stimuli. Significant advancements were achieved by incorporating multimodal data, such as text (Lin et al., 2022), and then by leveraging the generative power of latent diffusion models (LDMs) (Takagi and Nishimoto, 2023; Scotti et al., 2023; Lu et al., 2023; Xia et al., 2024). A more detailed literature review is presented in Appendix 2. However, these state-of-the-art methods require extremely large datasets with dozens of hours of fMRI data per participant. Thus, there is a clear need to maximize the efficiency and generalizability of image reconstruction methods so that a model for a new participant can be trained using as little data as possible.
Recent works by Scotti et al. (2024) and Liu et al. (2024) have attempted to improve generalization by projecting fMRI signals into a shared representation space across subjects. However, in these pipelines the shared space is treated only as a convergence point of subject-specific pathways and is not explicitly aligned. We show in our analysis that though subject representations in this space share a similar structure, manipulations such as a simple orthogonal transformation substantially increase cross-subject similarity. This motivates an explicit alignment strategy: instead of relying on the shared space to emerge implicitly, lightweight subject-specific neural network modules called adapters (Liu et al., 2024) can be trained to align each subject to a reference subject.
We propose Adapter Alignment (AA), an efficient and generalizable approach to fMRI-based visual reconstruction using a subject-agnostic representation space. Instead of fine-tuning new subjects end-to-end, we align new subjects to a pre-trained reference subject using adapter modules. AA substantially reduces data and training time while improving reconstruction quality. We also present a data-efficient image selection strategy that identifies the most representative training samples by covering key dimensions of variance in the common space. This method reduces the burden of data collection and further complements Adapter Alignment by making fine-tuning more efficient. Our contributions are as follows:
  • We demonstrate that fMRI signals from different subjects exhibit structural similarity in a shared representation space. While orthogonal transformations improve alignment, near-perfect cross-subject alignment can be attained using adapter modules (Section 3).
  • We introduce Adapter Alignment (AA), a novel training paradigm where new subjects are explicitly aligned to a reference subject. AA achieves faster convergence and better performance, especially in low-data regimes (Sections 4.2, 5.1 and 5.2).
  • We propose a submodular greedy algorithm to identify representative training images, providing a data-efficient alternative to random sampling. This strategy consistently improves fine-tuning performance across subjects and complements Adapter Alignment by reducing the amount of data required. (Sections 4.4 and 5.2)
  • We validate AA on both the NSD and THINGS  (Hebart et al., 2023) datasets, showing generalization across different architectures (Section 5.4), data acquisition protocols (7T vs. 3T) (Section 6), and subject pools (Appendix C).
2 Related Work
Recent works have achieved impressive results by mapping fMRI data to latent diffusion model (LDM) spaces  (Takagi and Nishimoto, 2023; Scotti et al., 2023; Lu et al., 2023; Xia et al., 2024), while simultaneously integrating multiple modalities. Despite this progress, these methods have not been thoroughly tested for their generalization performance across a larger population of subjects. In other words, constructing a model for a new subject still requires retraining the model from scratch to ensure optimal performance.
Approaches like Ferrante et al. (2024) have demonstrated cross-subject brain decoding using limited data samples by aligning multiple subjects. Although the authors, like us, leveraged visual stimuli commonly used for multiple subjects, the fidelity of their reconstructed images lags behind current advancements. Scotti et al. (2024), produced significant results in image reconstruction under limited data conditions but did not fully exploit the properties of the common representation space. In contrast, our novel approach enables faster convergence and improved performance in a limited data setting by effectively leveraging the structure of the common space.
Qian et al. (2023) refers to a ’common space’; however, their representation is primarily anatomical and only implicitly semantic. Thual et al. (2022) focuses on aligning the entire cortical surfaces of different individuals in a way that preserves both anatomical structure and functional activation patterns. However many downstream tasks like mapping fMRI signals to semantic embeddings for image reconstruction benefit from representations that capture high-level semantic content. Therefore the suggested approach doesn’t necessarily capture abstract, semantic features (such as object categories, scenes, or other high-level concepts) that are crucial for decoding complex stimuli. Thual et al. (2023) is an extension of Thual et al. (2022). They apply functional alignment on top of the anatomical alignment. The method is very efficient even in limited data settings but it’s less integrated, meaning that it appears more as a preprocessing step. Piecewise shared response modeling Bazeille et al. (2021) provides significant performance in terms of aligning neural data from different subjects. However, the independent alignment of segments might not integrate seamlessly into a coherent global representation. Similarly to MindEye2, but in the context of M/EEG data, Li et al. (2024) manage to reconstruct visual stimuli using cross subject decoding. They tackle inter-subject variability primarily by incorporating subject-specific tokens into their M/EEG encoder.
@@ 3 A Common Visual Representation Space
At the core of our approach is the hypothesis that each individual’s brain (or at least the tissue involved in processing static visual images) can be viewed as an instance of a shared ’common brain’, with visual activity patterns that vary along similar dimensions in a shared representational space. This idea is supported by prior neuroscience findings, including consistent category-selective tissue responses (e.g., faces, scenes) and continuous dimensions of high-level visual selectivity in ventral temporal cortex such as animate/inanimate, large/small (real-world size), and spiky/round (Huth et al., 2012; Long et al., 2018; Bao et al., 2020). Retinotopic organization provides another example, where adaptive models transform individual anatomical mappings from retinal to cortical coordinates into a shared reference space (Sprague and Serences, 2013; Wandell and Winawer, 2015). In our work, we analyze embeddings from multiple subjects and provide evidence that although fMRI signals do not naturally align in a shared space, they exhibit a similar underlying structure. We further explore structural and semantic properties of this space through eigenvector analysis, visual embedding comparisons, and alignment experiments, with additional results of semantic categories (e.g. animate / inanimate) provided in Appendix J.
@@
Existing works, such as MindEye2 (Scotti et al., 2024), employ simple linear adapters to project subject-specific fMRI signals into a common representation space. A key indication of the existence of such a space is precisely this structural similarity across subjects. To investigate it, we use a pre-trained MindEye2 model to extract common-space embeddings for shared images across multiple subjects. Initial cross-subject cosine similarity is low, but as shown in Figure 1, applying an orthogonal transformation with Procrustes analysis (Schönemann, 1966) reveals that the embeddings can be structurally aligned. However, this alignment is still incomplete, reflecting both the limitations of orthogonal transformations and the absence of an explicit alignment objective.
(a) Original alignment.
(b) Post orthogonal transformation.
(c) Post Adapter Alignment.
Figure 1: Cross-subject cosine similarity of shared image embeddings in the common space extracted from a pre-trained MindEye2 model. (a) Alignment between subjects is weak before any transformation. (b) Applying an orthogonal transformation substantially improves cross-subject similarity. (c) Explicit alignment through lightweight adapter modules achieves near-perfect alignment. Only Subjects 1, 2, 5, and 7 are shown for consistency throughout the manuscript.
As a preliminary experiment, we pre-trained a single-subject model and explicitly mapped subsequent subjects to this reference subject using lightweight subject-specific adapters. Even with a simple adapter design (a linear layer with a GELU), this approach yielded near-perfect alignment in the common space (Figure 1 (c)). The choice of adapter is not critical. Any transformation that goes beyond an orthogonal mapping can substantially improve alignment highlighting that it is the explicit alignment objective, rather than architectural complexity, that drives the improvement.
The above evidence of structurally well-aligned representation spaces across subjects forms the core hypothesis of our work. In the remainder of this paper, we develop a training paradigm that explicitly enforces such alignment and evaluate it empirically against the commonly used baseline fine-tuning approach. Additional analysis is provided in Appendix I.
4 Methods
In the previous section, we established the existence of a shared representation space across subjects. We now describe the methods used to leverage this space for fMRI-to-image reconstruction. Our framework introduces Adapter Alignment (AA) as a lightweight strategy for subject-specific alignment, integrates this into a reconstruction pipeline, and augments it with a greedy image selection algorithm to further improve data efficiency. Together, these components form the basis for the experiments presented in the following section.
@@ 4.1 Dataset
We use the Natural Scenes Dataset (NSD) (Allen et al., 2022) but apply a different split compared to the traditional setup. Typically, common images across subjects are assigned to the test set, while the training set consists of unique images viewed by each subject. However, for effective Adapter Alignment, it is critical to have a one-to-one mapping of images across subjects. To achieve this, we incorporate the common images into the training set and swap an equal amount of training images to the test sets. This means that every subject now has a unique set of images on which they are tested. The test split is chosen to roughly ensure an equal distribution across different categories in the dataset. We present the performance of the original training method and the proposed technique on the new split to ensure a fair comparison across subjects and alignment techniques. Three splits are generated in this manner and whenever results are presented on the new split, it is an average of the performance on all 3 splits.
@@
Figure 2: Training procedures for fMRI-Image Reconstruction. Top: Traditional Approach to reconstruction. All subjects are trained simultaneously and loss functions are applied at the MLP Mapper to align the output to pre-trained CLIP space. Bottom: Proposed AdapterAlign training pipeline. First, we train a single reference subject and obtain its embeddings in the shared space. Then we train subsequent subjects to minimize loss at both the shared space and the mapper. The same mapper weights are transferred over in step 2.
@@ 4.2 Adapter Alignment
Current approaches for training multi-subject pipelines with a common representation space typically follow an end-to-end strategy, where fMRI inputs are processed by the model and the loss is computed at the final layer to minimize the discrepancy between the ground-truth and predicted embeddings. Throughout the paper, we refer to this fine-tuning strategy as the baseline. In this setup, the common space emerges organically during training without explicit constraints. However, this approach has two key drawbacks: (1) it requires extended training to form a satisfactory common space and achieve convergence, and (2) the subject-specific embeddings in the common space are not strongly aligned, making it harder for the residual MLP to map multiple subject inputs to a unified output (Figure 2).
@@
To address these challenges, we propose Adapter Alignment (AA) training. In AA training, the pipeline is first pre-trained end-to-end on a single reference subject, chosen based on reconstruction performance and data availability. Among NSD subjects, Subject 1 consistently exhibited the best reconstruction performance, making it the ideal candidate to construct the reference space. After pre-training, the output embeddings for the subject-specific adapter, particularly those corresponding to the shared images, are extracted.
@@
When fine-tuning a new subject, we adopt a two-step process. The first is Adapter Alignment, where we train the new subject’s adapter using the shared images, with the objective of aligning its output to the embeddings of Subject 1 for the same images. The adapter is allowed to overfit on these shared images until the loss is minimized. The second step is End-to-End Fine-tuning, where we unfreeze the entire pipeline and resume training, applying the usual losses at the final output. Additionally, we introduce an MSE loss at the adapter level specifically for the common images, which allows the adapter to restructure itself to account for subject-specific discrepancies and adapt to the unique images encountered by the new subject. We refer to this process of overfitting the adapter with AA and fine-tuning end-to-end "AAMax". From our preliminary experiments, we found that AAMax outperforms other methods of aligning adapters like the ones proposed by Ferrante et al. (2024) and Wang et al. (2024) Both methods do not fine-tune the MLP and the diffusion prior end-to-end after aligning the adapters which reduces the generalization ability of the fine-tuned models. We compare the performance between these training methods in Appendix E. We find that AAMax is the best method for fine-tuning new subjects in limited data settings.
@@
This training method is effective across both high-data and low-data regimes. With full training data, AAMax achieves performance comparable to traditional end-to-end training while converging more quickly. In limited data settings, AAMax shows a clear advantage, producing higher-quality reconstructions and greater accuracy with substantially fewer training samples. We outline our method in Figure 2 and detail each component of the pipeline in the following section.